# Distributionally Robust Markov Games with Average Reward

### v2

@@ Appendix F Proof of Section 7
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted distributionally robust Markov game by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward distributionally robust Markov game.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted distributionally robust Markov game is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
@@
due to $\pi$ is an $\epsilon$-NE for the discounted robust MG, which hence completes the proof. ∎
Lemma F.1.
The discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2) of Theorem 7.1.
Proof.
The proof is a direct corollary of Theorem 7.1 and results from (Roch2025reduction). Specifically, it is shown in Theorem 3.4 in (Roch2025reduction) that, for any $\gamma>1-\frac{\epsilon}{D}$ [^2] The results in (Roch2025reduction) is derived for $\mathcal{H}$ which is the robust optimal span. However, their results also hold for any constant that is larger than $\mathcal{H}$. Our claim thus holds by noting that $D\geq\mathcal{H}$ (Roch2025reduction)., it holds that

### v3

@@ Appendix F Proof of Section 7
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted DR-MG by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward DR-MG.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted DR-MG is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
(3). The discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2)
@@
due to $\pi$ is an $\epsilon$-NE for the discounted robust MG, which hence completes the proof.
(3). The proof is a direct corollary of Theorem 7.1 and results from (Roch2025reduction). Specifically, it is shown in Theorem 3.4 in (Roch2025reduction) that, for any $\gamma>1-\frac{\epsilon}{D}$ [^3] The results in (Roch2025reduction) is derived for $\mathcal{H}$ which is the robust optimal span. However, their results also hold for any constant that is larger than $\mathcal{H}$. Our claim thus holds by noting that $D\geq\mathcal{H}$ (Roch2025reduction)., it holds that